# WASUP: Interpretable Classification with Weight-Input Alignment and Class-Discriminative SUPports Vectors

### v1

@@ 3.1 Architecture
Figure 3: WASUP consists of a feature extractor $\mathcal{F}$ with model weights $\theta$ that extracts latent vectors $f$. The evidence predictor $\mathcal{E}$ converts them into positive-valued vectors $f^{+}$, which, given other images from the training set, can also serve as support vectors $v_{i}^{c}$. Then, it computes the similarity between the input image’s latent vector $f$ and support vectors $v_{i}^{c}$. Summing the class-specific temperature-normalized similarity scores and adding a bias $b$ yields the class logit $\mu_{c}$. The B-cos transform facilitates the summary of a forward pass as a single weight-matrix $\mathbf{W}_{1\rightarrow L}(\mathcal{I})$, which is leveraged to compute faithful contribution maps.
WASUP consists of a feature extractor $\mathcal{F}_{\theta}$ (see Fig. 3) employs a B-cos backbone and transforms an image $\mathcal{I}$ into latent vectors $f=\mathcal{F}_{\theta}(\mathcal{I})\in\mathbb{R}^{d}$, with $\mathcal{I}\in\mathbb{R}^{ch\times H\times W}$, with $d$ being the latent dimension, $ch$, $H$, and $W$ the image’s channels, height, and width, respectively, and $\theta$ the model weights ($\mathcal{F}_{\theta}$ can be any B-cos neural network architecture).
The evidence predictor $\mathcal{E}$ consists of a function $\oplus:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}_{\geq 0}$ that transforms the real-valued vectors $f$ into non-negative vectors $f^{+}=\oplus(f)$, and a similarity measure that computes similarities between support vectors $v_{i}^{c}\in F^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}$ and $f^{+}$ as $sim(f^{+},v_{i}^{c})\in\mathbb{R}$, with $\Omega(F^{c})=N_{s}$ and $i$ denoting the $i$-th support vector of class $c$. Logits are computed for class $c$ as the sum of its similarity scores:
$\mu_{c}=b+\sum_{v_{i}^{c}}\frac{sim(f^{+},v_{i}^{c})}{\mathcal{T}},$ (5)
where $b$ represents a fixed bias term, $i=1,\dots,N_{s}$ indexes the support samples, and $\mathcal{T}$ is a temperature to improve convergence by scaling the logits’ magnitude appropriately. Examples of suitable transformation functions $\oplus$ include the exponential function, absolute value, and ReLU.
@@ 3.2 Training and Optimization
Following Wang & Sabuncu (2022), we randomly sample the training set to extract $N_{s}$ support vectors $v_{k}^{c}$ per class to encourage the model to form distinct, class-specific clusters in latent space. After every epoch, we compute sets of latent vectors $\hat{V}^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}_{k=1}^{N}$, which we each cluster using k-means to extract centroids $\gamma_{i}^{c}$. Then, we replace each centroid $\gamma_{i}^{c}$ with the closest latent vector to yield support vectors from the same class, defined as:
$v_{i}^{c}=\arg\min_{f^{+}_{k}\mid y_{k}=c}\|f^{+}_{k}-\gamma_{i}^{c}\|_{2},$ (6)
which facilitates case-based reasoning similar to k-nearest-neighbors, as only features descriptive of the class can increase the probability of that class by definition of WASUP.
We follow the argument of Böhle et al. (2024) that training with binary cross entropy loss increases the alignment pressure induced by B-cos networks, and compute the loss for every batch of $N_{b}$ samples and $C$ classes as:
$\mathcal{L}_{BCE}=-\frac{1}{N_{b}}\sum_{\hat{b}=1}^{N_{b}}\sum_{c=1}^{C}\left[y_{\hat{b},c}\log(\hat{y}_{\hat{b},c})+(1-y_{\hat{b},c})\log(1-\hat{y}_{\hat{b},c})\right],$ (7)
with $\hat{y}_{\hat{b},c}=\sigma(\mu_{\hat{b},c})$, $\sigma$ denoting the sigmoid function. If trained with cross-entropy loss, it could potentially converge towards optima in which activation of the wrong classes is still relatively large, and thus, the alignment pressure would fail.
@@ 3.3 Explainability
While global explanations provided by WASUP are ideal for a developer to debug a model, they also facilitate case-based reasoning, which is most intuitive to understand for humans (Nguyen et al., 2021; Kim et al., 2023; Jeyakumar et al., 2020).
@@
Additionally, WASUP’s decision-making is based on a fixed set of latent support vectors. Hence, we can compute the inter- and intra-class similarities to gain insights into the latent space learned by the feature extractor.
@@
Local explanations, shown in Fig. 2 are threefold: First, the temperature-scaled similarity scores $\frac{sim(f^{+},v_{i}^{c})}{\mathcal{T}}$ (support evidence) provided by the evidence predictor $\mathcal{E}$ quantify the alignment of a test image with a support sample, and their proportion of the log-probability. Second, since a test image is assigned to a class if it is similar to that class’s support sample (see Fig. 2), understanding which parts of the support sample’s image are compressed in the support vectors themselves is crucial, i.e., explaining the model mechanisms. The B-cos properties of WASUP allow us to faithfully compute the RGBA explanations in terms of the similarity measure. Since each prediction can be summarized by a linear equation, we compute contribution maps in terms of the class logit (test contribution) following the linearity axiom (Sundararajan et al., 2017) ($\phi=\alpha\phi_{1}+\beta\phi_{2}\text{ iff }\mathcal{F}(x)=\alpha\mathcal{F}_{1}(x)+\beta\mathcal{F}_{2}(x)$). Therefore, the test sample contributions highlight the image features that aligned with the support vectors. Third, the global RGBA explanations (support contributions) of the support samples allow the user to check for the intersection of support contributions and test contributions, similar to k-nearest-neighbor.
@@
To thoroughly assess the interpretability of WASUP, we conduct a theoretical evaluation of its explanations based on the axioms defined by Sundararajan et al. (2017), namely Completeness, Sensitivity, Implementation Invariance, Dummy, Linearity, and Symmetry-Preserving. Our proofs (see App. A.1), confirm that WASUP’s explanations adhere to these principles, thereby validating the model’s capability to provide faithful and comprehensive insights into its decision-making processes.
@@ 4 Experiments and Results
We evaluate WASUP on three tasks with three backbone architectures as feature extractors. The tasks are (i) multi-label image classification on Pascal VOC (Everingham et al., 2009), (ii) fine-grained image classification with mutually exclusive labels on Stanford Dogs (Aditya et al., 2011), (iii) pathology prediction with mutually exclusive labels on RSNA (Shih et al., 2019). The backbone architectures are B-cos versions of DenseNet121 (Huang et al., 2017), ResNet50 (He et al., 2016), and a Hybrid Vision Transformer (Xiao et al., 2021) (VitC), with $\sim$8M, $\sim$26M, and $\sim$81M trainable parameters, respectively.
Compared to black-box versions of the backbones (see Tab. A.2), WASUP performs on par on Pascal VOC ($\geq+0.29\%$), worse on Stanford Dogs ($-3.7\%$ Acc), and comparable on RSNA ($-0.65\%$ Acc); take into account that we train the model from scratch on RSNA.
@@
Pascal VOC (Everingham et al., 2009)
Fig. 4 illustrates explanations of WASUP for multi-label image classification on Pascal VOC with a DenseNet121 backbone. As seen in Figs. 4 (a) and (b), WASUP shifts its focus in the test image depending on the class to predict, i.e., the corresponding column $c$ in $\left[\mathbf{W}_{1\rightarrow L}\right]^{T}_{c}$: To predict the cat class, only the area around the cat is present in the contribution map. When predicting the human class, for which the model computed very high support evidence, the complementary area of the image is contained in the contribution map. The support contributions show that the model is particularly sensitive to the eyes of a cat. In the test sample in Fig. 4 (a) and (b), only the left half and one eye of the cat are visible, which explains the comparably low support evidences in this particular instance. Arguably, the support contributions of the human class in Fig. 4 (b) suggest that the model learned to look for human eyes in support sample 42, which shows high contributions from the two lights on the ceiling. For support samples 43 and 44, the whole heads of the humans are the primary contributions to the support vectors.
@@
We present additional results for other test images and backbones in Figs. A.5-A.7. In many test sample instances (e.g., compare Fig. A.5 (a) vs. (b), Fig. A.5 (c) vs. (d), Fig.A.7 (a) vs. (b)), we observe a strong shift in contributions depending on the class prediction. This demonstrates that the latent vectors of samples with multiple labels efficiently encode features for all classes that need to align with the class support vectors. In Fig. A.6 (a)-(d), we see that the model predicted four classes, for which, according to the dataset labels, only a single class was actually contained (chair). However, the explanations look very reasonable. Indeed, there is a table in the scene; whether it is a dining table is obviously debatable. There is also a child in the image, so predicting humans is not unreasonable. Lastly, the potted plant class logit is very low, and the test contribution map highlights the flower on the table.
@@
For the test sample predicted in Fig. A.7 (d) and (e), we see the correct classification of humans with high support evidence scores (note how the model extracted the human for support sample 44 in very dark lighting conditions), and WASUP believed that the martial arts club logo on the person’s outfit resembled the logo of a drink.
Stanford Dogs (Aditya et al., 2011)
@@
We evaluate WASUP with a ResNet50 backbone on the Stanford dogs dataset. The contribution maps for the test samples in Figs. A.8,A.9,A.10 show that WASUP consistently focuses on the dogs and discards background elements such as humans, especially in Figs. A.8 (a) and (b).
@@
The wrong classification of the test sample presented in Fig. A.8 (c) shows that primarily the head of the dog contributed to an increased similarity with the support vector of support sample 46, whose support contribution focuses on the head, too. In contrast, the support contributions of support samples 45 and 47 both indicate that the support vectors contain large portions of the body of the dog. Evaluating the explanations of the correct class in Fig. A.8 (d) shows that the support sample 33 containing two dogs has high support evidence, as there are two dogs of its class in the test image. However, the low activation of support samples 34 and 35, which both contain one dog, resulted in little support evidence, indicating that the number of instances is encoded into support vectors.
@@
The different magnitude in the support evidence suggests that the model benefits from increasing the number of support samples in settings with high intra-class variability, which manifests in the visualization of support samples and their contribution maps. Considering the support evidence of support sample 162 in Figs. A.10 (a) and (b), its support evidence for the test sample in Fig. A.10 (a) is lower compared to the test sample in Fig. A.10 (b), which only contains the lower portion of the head like support sample 162. The support contributions of the other support samples that focus on the eyes and the body show the inverse behavior of the two respective test samples. For some classes, we observe similar support evidence for all support samples, e.g. Fig. A.9 (a) and (b), suggesting lower intra-class variability within the training data for the particular classes.
RSNA (Shih et al., 2019)
@@
We utilize the RSNA dataset to showcase WASUP on a decision-critical task to differentiate between images with lung opacity and others (healthy and unhealthy without lung opacity, indicated by the presence of bounding boxes) from chest X-ray images. Here, we want to verify reasonable decision-making by evaluating explanations in terms of bounding boxes. An explanation of a classifier trained to predict lung opacity should focus on an area inside the bounding box, illustrated in Figs. A.11-A.12.
@@
The predictions in Figs. A.11 (a)-(d) exemplify how the support evidence varies across the class’s support samples and suggest that the increasing number of support samples allowed WASUP to reproduce the class’s latent space with a likely high variability. Additionally, we observe a high degree of intersection between significant test sample contributions and bounding boxes for the occluded class in Figs. A.11 (a)-(d) and that the focus of the model is within the lungs for non-occluded samples in Fig. A.12, indicating that the model indeed learned to differentiate occlusion from no-occlusion using medically reasonable regions.
Debugging WASUP
@@
While the experiments primarily highlighted the explainability features from the view of a user, we now exemplify how the global and local explanations help to identify flaws in the optimization process. Revisiting the model trained on RSNA, we detail how to use the global explanations of WASUP to gain more insights into the model behavior. We observe that the support evidence of support sample 0 of the correct classifications in Figs. A.12 (a) and (b) is almost absent compared to the wrong classification in Fig. A.12 (c). Considering that support contribution 0 contains large parts of the input, this raises suspicion. In addition, we see that the only significant evidence for the correct class depicted in Fig. A.12 (d) comes from support sample 0, which suffers from contributions across the whole input as well. As a result, we evaluate the latent space learned by WASUP in Fig. A.14 and find that support vectors 0 and 3, although of different classes, are very close in the t-SNE projection. We attribute this to a weak latent representation, suggesting that the model has yet to learn a more class-separable latent space.
@@
Next, we evaluate whether the model indeed learned class-representative support vectors (see Fig. A.15) by computing the Silhouette score (0.655) (Rousseeuw, 1987) and plotting the inter- and intra-class B-cos similarities of WASUP trained on Pascal VOC with a DenseNet121 backbone in Fig. A.16. We observe that the intra-class similarities on the diagonal are indeed higher than those off-diagonal, suggesting good class distinction in latent space. Additionally, the average distance between some classes is a bit higher, e.g., dining table, chair, and bottle. This is expected, as images containing a dining table often include a chair and a bottle, requiring the model to extract a latent vector capable of predicting all three classes with sufficient confidence.

### v2

@@ 3.1 Architecture
Figure 1: SIC consists of a feature extractor $\mathcal{F}$ with model weights $\theta$ that extracts latent vectors $f$. The evidence predictor $\mathcal{E}$ converts them into positive-valued vectors $f^{+}$, which, given other images from the training set, can also serve as support vectors $v_{i}^{c}$. Then, it computes the similarity between the input image’s latent vector $f$ and support vectors $v_{i}^{c}$. Summing the class-specific temperature-normalized similarity scores and adding a bias $b$ yields the class logit $\mu_{c}$. The B-cos transform facilitates the summary of a forward pass as a single weight-matrix $\mathbf{W}_{1\rightarrow L}(\mathcal{I})$, which is leveraged to compute faithful contribution maps.
We visualize SIC in Fig. 1. It consists of a B-cos feature extractor $\mathcal{F}_{\theta}$, which transforms an image $\mathcal{I}$ into latent vectors $f=\mathcal{F}_{\theta}(\mathcal{I})\in\mathbb{R}^{d}$, with $\mathcal{I}\in\mathbb{R}^{ch\times H\times W}$, $d$ the latent dimension, $ch$, $H$, and $W$ the image’s channels, height, and width, respectively, and $\theta$ the model weights ($\mathcal{F}_{\theta}$ can be any B-cos neural network architecture).
The evidence predictor $\mathcal{E}$ consists of a function $\oplus:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}_{\geq 0}$ that transforms the real-valued vectors $f$ into non-negative vectors $f^{+}=\oplus(f)$, and a similarity measure that computes similarities between support vectors $v_{i}^{c}\in F^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}$ and $f^{+}$ as $sim(f^{+},v_{i}^{c})\in\mathbb{R}$, with $\Omega(F^{c})=N_{s}$ and $i$ denoting the $i$-th support vector of class $c$. Logits are computed for class $c$ as the sum of its similarity scores $\mu_{c}=b+\sum_{v_{i}^{c}}\frac{sim(f^{+},v_{i}^{c})}{\mathcal{T}}$, where $b$ represents a fixed bias term, $i=1,\dots,N_{s}$ indexes the support samples, and $\mathcal{T}$ is a temperature to improve convergence by scaling the logits’ magnitude appropriately. Examples of suitable transformation functions $\oplus$ include the exponential function, absolute value, and ReLU.
@@ 3.2 Training and Optimization
Following Wang and Sabuncu [31], we randomly sample the training set to extract $N_{s}$ support vectors $v_{k}^{c}$ per class to encourage the model to form distinct, class-specific clusters in latent space. After every epoch, we compute sets of latent vectors $\hat{V}^{c}=\{f^{+}_{k}=\mathcal{F}_{\theta}(\mathcal{I}_{k})~{}|~{}y_{k}=c\}_{k=1}^{N}$, which we each cluster using k-means to extract centroids $\gamma_{i}^{c}$. Then, we replace each centroid $\gamma_{i}^{c}$ with the closest latent vector to yield support vectors from the same class, defined as $v_{i}^{c}=\arg\min_{f^{+}_{k}\mid y_{k}=c}\|f^{+}_{k}-\gamma_{i}^{c}\|_{2}$, which facilitates case-based reasoning similar to k-nearest-neighbors, as only features descriptive of the class can increase the probability of that class by definition of SIC.
We follow Böhle et al. [7] and optimize the model with binary cross-entropy loss, which effectively increases alignment pressure compared to training with standard multi-class cross-entropy loss.
@@ 3.3 Explainability
While global explanations provided by SIC are ideal for a developer to debug a model, they also facilitate case-based reasoning, which is most intuitive to understand for humans [22, 17, 15].
@@
Additionally, SIC’s decision-making is based on a fixed set of latent support vectors. Hence, we can compute the inter- and intra-class similarities to gain insights into the latent space learned by the feature extractor.
@@
Local explanations are threefold: First, the temperature-scaled similarity scores $\frac{sim(f^{+},v_{i}^{c})}{\mathcal{T}}$ (support evidence) provided by the evidence predictor $\mathcal{E}$ quantify the alignment of a test image with a support sample, and their proportion of the log-probability. Second, since a test image is assigned to a class if it is similar to that class’s support sample, understanding which parts of the support sample’s image are compressed in the support vectors themselves is crucial, i.e., explaining the model mechanisms. The B-cos properties of SIC allow us to faithfully compute the RGBA explanations in terms of the similarity measure. Since each prediction can be summarized by a linear equation, we compute contribution maps in terms of the class logit (test contribution) following the linearity axiom [30] ($\phi=\alpha\phi_{1}+\beta\phi_{2}\text{ iff }\mathcal{F}(x)=\alpha\mathcal{F}_{1}(x)+\beta\mathcal{F}_{2}(x)$). Therefore, the test sample contributions highlight the image features that aligned with the support vectors. Third, the global RGBA explanations (support contributions) of the support samples allow the user to check for the intersection of support contributions and test contributions, similar to k-nearest-neighbor.
@@
To thoroughly assess the interpretability of SIC, we conduct a theoretical evaluation of its explanations based on the axioms defined by Sundararajan et al. [30], namely Completeness, Sensitivity, Implementation Invariance, Dummy, Linearity, and Symmetry-Preserving. Our proofs (see App. A.1), confirm that SIC’s explanations adhere to these principles, thereby validating the model’s capability to provide faithful and comprehensive insights into its decision-making processes. Thus, SIC solves the limitations of previous work, which either does not provide pixel-level explanations, like the NW-Head [31] or lack global explanations like B-cos neural networks [7].
@@ 4 Experiments and Results
We evaluate SIC on three tasks with three backbone architectures as feature extractors. The tasks are (i) multi-label image classification on Pascal VOC [10], (ii) fine-grained image classification with mutually exclusive labels on Stanford Dogs [2], and (iii) pathology prediction with mutually exclusive labels on RSNA [29]. Additionally, we evaluate the explanations given by SIC on the FunnyBirds framework [12]. The backbone architectures are B-cos versions of DenseNet121 [14], ResNet50 [11], and a Hybrid Vision Transformer [34] (VitC), with $\sim$8M, $\sim$26M, and $\sim$81M trainable parameters, respectively.
First, we compare the performance of SIC in terms of the (balanced) accuracy to black-box neural networks and established inherently interpretable neural networks, namely ProtoPNet [8], a BagNet17 [6], the NW-Head [31], and B-Cos neural networks [7] in Fig. 2 and Tab. A.2. Compared to black-box versions of the backbones, SIC performs on par on Pascal VOC and RSNA ($\geq+0.29\%$ and $-0.65\%$ Acc, respectively), and worse on Stanford Dogs ($-3.7\%$ Acc). Overall, ProtoPNet and BagNets perform worse than all competing methods, and SIC performs better than nine out of 14 models.
Figure 2: Results in terms of accuracy for ResNet backbones on Pascal VOC and Stanford Dogs, and DenseNet backbones on RSNA. SIC performs superior to nine out of 14 models. Refer to Tab. A.2) for exhaustive results.
@@
Pascal VOC [10]
Fig. 3 illustrates explanations of SIC for multi-label image classification on Pascal VOC with a DenseNet121 backbone. As seen in Figs. 3 (a) and (b), SIC shifts its focus in the test image depending on the class to predict, i.e., the corresponding column $c$ in $\left[\mathbf{W}_{1\rightarrow L}\right]^{T}_{c}$: To predict the cat class, only the area around the cat is present in the contribution map. When predicting the human class, the complementary area of the image is contained in the contribution map. The support contributions show that the model is particularly sensitive to the eyes of a cat. In the test sample in Fig. 3 (a) and (b), only the left half and one eye of the cat are visible, which explains the comparably low support evidences in this particular instance. Arguably, the support contributions of the human class in Fig. 3 (b) suggest that the model learned to look for human eyes in support sample 42, which shows high contributions from the two lights on the ceiling. For support samples 43 and 44, the whole heads of the humans are the primary contributions to the support vectors.
@@
We present additional results for other test images and backbones in Figs. A.7-A.9. In many test sample instances (e.g., compare Fig. A.7 (a) vs. (b), Fig. A.7 (c) vs. (d), Fig.A.9 (a) vs. (b)), we observe a strong shift in contributions depending on the class prediction. This demonstrates that the latent vectors of samples with multiple labels efficiently encode features for all classes that need to align with the class support vectors. In Fig. A.8 (a)-(d), we see that the model predicted four classes, for which, according to the dataset labels, only a single class was actually contained (chair). However, the explanations look very reasonable. Indeed, there is a table in the scene; whether it is a dining table is obviously debatable. There is also a child in the image, so predicting humans is not unreasonable. Lastly, the potted plant class logit is very low, and the test contribution map highlights the flower on the table.
@@
For the test sample predicted in Fig. A.9 (d) and (e), we see the correct classification of humans with high support evidence scores (note how the model extracted the human for support sample 44 in very dark lighting conditions), and SIC believed that the martial arts club logo on the person’s outfit resembled the logo of a drink.
Stanford Dogs [2]
@@
We evaluate SIC with a ResNet50 backbone on the Stanford dogs dataset. The contribution maps for the test samples in Figs. A.10A.11,and  A.12 show that SIC consistently focuses on the dogs and discards background elements such as humans, especially in Figs. A.10 (a) and (b).
@@
The wrong classification of the test sample presented in Fig. A.10 (c) shows that primarily the head of the dog contributed to an increased similarity with the support vector of support sample 46, whose support contribution focuses on the head, too. In contrast, the support contributions of support samples 45 and 47 both indicate that the support vectors contain significant portions of the body of the dog. Evaluating the explanations of the correct class in Fig. A.10 (d) shows that the support sample 33 containing two dogs has high support evidence, as there are two dogs of its class in the test image. However, the lower support evidence of 34 and 35, which both contain one dog, indicate that the number of instances is encoded into support vector 33.
@@
Considering the support evidence of support sample 162 in Figs. A.12 (a) and (b), its support evidence for the test sample in Fig. A.12 (a) is lower compared to the test sample in Fig. A.12 (b), which only contains the lower portion of the head like support sample 162. The support contributions of the other support samples that focus on the eyes and the body show the inverse behavior of the two respective test samples. However, we observe alike support evidence for all support samples for other classes, e.g., Fig. A.11 (a) and (b). This exemplifies that SIC is able to capture the distribution of more challenging classes.
RSNA [29]
@@
We utilize the RSNA dataset to showcase SIC on a decision-critical task to differentiate between images with lung opacity and others (healthy and unhealthy without lung opacity, indicated by the presence of bounding boxes) from chest X-ray images. Here, we want to verify reasonable decision-making by evaluating explanations in terms of bounding boxes. An explanation of a classifier trained to predict lung opacity should focus on an area inside the bounding box, illustrated in Figs. A.13 and A.14.
@@
The predictions in Figs. A.13 (a)-(d) exemplify how the support evidence varies across the class’s support samples and suggest that the increasing number of support samples allowed SIC to reproduce the class’s latent space with a likely high variability. Additionally, we observe a high degree of intersection between significant test sample contributions and bounding boxes for the occluded class in Figs. A.13 (a)-(d) and that the focus of the model is within the lungs for non-occluded samples in Fig. A.14, indicating that the model indeed learned to differentiate occlusion from no-occlusion using medically reasonable regions.
Figure 4: Results on the FunnyBirds Framework [12]. Note that Completeness (Com.) is the mean across Controlled Synthetic Data Check (CSDC), Preservation Check (PC), Deletion Check (DC), and Distractability (D); Accuracy (A), Background Independence (BI), Correctness(Cor.), Contrastivity (Con.). Refer to Tab. A.4 for full results.
FunnyBirds [12]
This framework contains a dataset of synthetic renderings of 50 bird species that are defined by the five concepts beak, wings, feet, eyes, and tail. Removing or replacing a concept with one of another class enables analysis of explanations in terms of completeness, background independence, contrastivitiy, and correctness.
The results, presented in Figs. 4 and A.5, and Tab. A.4, show that SIC outperforms ProtoPNet [8] and B-cos [5] in eight out of nine explainability metrics. However, BagNet [6] has a lead in terms of accuracy (+1%), correctness (+4%), and contrastivity (+4%). The explanations for the representative test sample in Fig. A.6 show that SIC extracted a diverse range of poses with each support sample offering a distinct view of the relevant concepts. Moreover, SIC demonstrates near-perfect resilience to background noise, evidenced by a background independence of 0.99 and the support contributions shown in Fig. A.6. A contrastivity of 0.95 further confirms that only class-discriminative features were learned, which aligns with our model design intent. Correctness reflects the assumption that iterative removal of the most important parts from a test image produces a corresponding decline in the classification probability. SIC achieves 0.72 correctness, which is in line with the scores ($<0.80$) reported by Hesse et al. [12]. A potential explanation is that the part importance is computed by summing attributions scores within a part without normalization for the part’s size. Finally, a completeness of 0.97 suggests that SIC encoded all class-representative features within its support vectors, as outlined in the support contributions in Fig. A.6.
Debugging SIC
@@
While the experiments primarily highlighted the explainability features from the view of a user, we now exemplify how the global and local explanations help to identify flaws in the optimization process. Revisiting the model trained on RSNA, we detail how to use the global explanations of SIC to gain more insights into the model behavior. We observe that the support evidence of support sample 0 of the correct classifications in Figs. A.14 (a) and (b) is almost absent compared to the wrong classification in Fig. A.14 (c). Considering that support contribution 0 contains large parts of the input, this raises suspicion. In addition, we see that the only significant evidence for the correct class depicted in Fig. A.14 (d) comes from support sample 0, which suffers from contributions across the whole input as well. As a result, we evaluate the latent space learned by SIC in Fig. A.16 and find that support vectors 0 and 3, although of different classes, are very close in the t-SNE projection. We attribute this to a weak latent representation, suggesting that the model has yet to learn a more class-separable latent space.
@@
Next, we evaluate whether the model indeed learned class-representative support vectors (see Fig. A.17) by computing the Silhouette score (0.655) [25] and plotting the inter- and intra-class B-cos similarities of SIC trained on Pascal VOC with a DenseNet121 backbone in Fig. A.18. We observe that the intra-class similarities on the diagonal are indeed higher than those off-diagonal, suggesting good class distinction in latent space. Additionally, the average distance between some classes is a bit higher, e.g., dining table, chair, and bottle. This is expected, as images containing a dining table often include a chair and a bottle, requiring the model to extract a latent vector capable of predicting all three classes with sufficient confidence.